OpenStudy (anonymous):

An usher passes out 60 percent of his programs before the intermission and 40 percent of the remainder after the intermission. At the end of the evening, what percent of the original number of programs does the usher have left? This is a remainder word problem which is why I can't make sense out of it. If the usher has 40 percent left, then he has nada left at the end right? But, the answer is 24. Could you please explain?

OpenStudy (anonymous):

I think what they're saying is that he hands out 60% of the programs to start with. He then hands out 40% of what he has left. So it's essentially asking what 100%-60% is, which is 40%, then it asks what 40% of that is, which would be 24

OpenStudy (anonymous):

I don't understand this part: which is 40%, then it asks what 40% of that is, which would be 24

OpenStudy (anonymous):

I think it's a little confusing, too. I'm sorry, I don't really know how to explain it.

OpenStudy (anonymous):

well, thanks =) i appreciate your help!

OpenStudy (anonymous):

Lets suppose he is writing 100 lines of program he wrote 60 lines before intermission and now he left out with 40 lines right And then he says 40% of the left program that is (40 Lines) is done after the intermission which means its not complete 40% of the program but it is 40% of the leftout program that is 40%of 40%

OpenStudy (anonymous):

yeah but that's 16 right?

OpenStudy (anonymous):

I got 16 when I did it

OpenStudy (anonymous):

It should be 40(60/100)

OpenStudy (anonymous):

I'm sorry I don't understand why.

OpenStudy (anonymous):

It says 40% of the programs left out so I assume it is 40 (100-40)% = 40(60%) = 40(60/100)